Please Factor the following 4 problems

21X^4 + 18X^2 + 36X^3
3T^2 - 5T - 2T^4
-3A^10 + 9A^5 - 6A^15
9M^3 - 7M^4 + 8M^2

Thank You for your help

Answer by WannabeCAgirl83(35)


21X^4 + 18X^2 + 36X^3 = x²(21x² + 18 + 36x)

3T^2 - 5T - 2T^4 = t(3t - 5 - 2t³)

-3A^10 + 9A^5 - 6A^15 = a^5(-3a^5 + 9 - 6a^10)

9M^3 - 7M^4 + 8M^2 = m²(9m - 7m² + 8)